OpenStudy (anonymous):

calculate lim x^2-1/x^2-x as x approahces to 1.

jimthompson5910 (jim_thompson5910):

(x^2-1)/(x^2-x) ((x+1)(x-1))/(x^2-x) ((x+1)(x-1))/(x(x-1)) (x+1)/x Now plug in x = 1 to get (x+1)/x (1+1)/1 2/1 2 So the answer is 2

OpenStudy (anonymous):

thanks.

OpenStudy (jamesj):

First notice that \[(x^2-1)/(x^2 - x) = (x-1)(x+1)/x.(x-1) = (x+1)/x\] Hence the limit \[ \lim_{x -> 1} (x^2-1)/(x^2 - x) = \lim_{x->1} (x+1)/x\ = 2/1 =2.\]

OpenStudy (anonymous):

thanks